OpenStudy (anonymous):

Solve the equation Quantity x plus 1 all over 3 = 2 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

(x+1)/3 = 2*x if this is what youre asking then the 3 which is for division would go up and multiply on the right hand side. Then x+1 = 6x. Now as the x on the left hand side moves to the right, it becomes negative, making 6x-x = 5x. Youre left with 1 on the RHS 1 = 5x therefore it becomes x = 1/5 (division of 1 by 5)

OpenStudy (anonymous):

So its x=5

OpenStudy (anonymous):

nono... x = 1/5.... 5*x = 1... So 5 goes down doesnt it :)

OpenStudy (anonymous):

Yea..well thx